OpenStudy (anonymous):

The sum of two consecutive integers is 59. Write an equation that models this situation and find the values of the two integers. A.) n+n+1=59;n=30;n+1=31 B.) n+2n=59;n=19;2n=38 C.) n+n+1=59;n=29;n+1=28 D.) n+n+1=59;n=29;n+1=30

OpenStudy (shane_b):

I recommend looking at each answer and trying to rule it out...

OpenStudy (anonymous):

and how do you do that? @Shane_B

OpenStudy (shane_b):

Let's take A for example. It states 3 things: 1) n+n+1=59 The "sum of two consecutive integers" is the same as saying n+n+1...and it's saying that it equals 59...so that looks ok. 2) n = 30 Maybe it doesn...don't know yet 3) n + 1 = 31 If n = 30 (as #2 says above) then n +1 = 31...so that looks ok. Now let's see if it all makes sense when you plug n=30 into the first equation (#1 above): 30 + 31 = 59 Nope...that's not right...so this answer cannot be correct

OpenStudy (shane_b):

If you do the same process for the other two then you will see which answer is correct

OpenStudy (anonymous):

i did the steps but im getting c but im not sure @Shane_B

OpenStudy (shane_b):

C states that n=29;n+1=28 If you look at that closely, you'll see that if n = 29, then n + 1 cannot equal 28

OpenStudy (anonymous):

ooo i see now its D @Shane_B

OpenStudy (shane_b):

You win :)

OpenStudy (anonymous):

yaayyyy thank you soo much @Shane_B

OpenStudy (shane_b):

No problem :)